# Vista: A Generalizable Driving World Model with High Fidelity and Versatile Controllability

### v3

@@ 2 Preliminary
Despite the high aesthetic quality, SVD lacks several key properties to function as a driving world model. As shown in Sec. 4, the prediction of SVD does not commence from the condition image, making it misaligned at the beginning and impractical for long-term extension. In addition, SVD struggles with the intricate dynamics of driving scenarios, entailing implausible motions. Moreover, SVD cannot be controlled by any action format. To this end, we aim to build a generalizable driving world model that predicts high-fidelity futures with realistic dynamics. It ought to be continuously extendable to long horizons and flexibly controllable by multi-modal actions as illustrated in Fig. 2.
@@ 3.2 Phase Two: Learning Versatile Action Controllability
Since different action formats cannot be converted into one another, it is impractical to apply multiple equivalent action conditions simultaneously at inference time. Additionally, it is burdensome to attain hybrid action controllability, which entails prohibitively expensive training to encompass all possible combinations. Thus, unlike common practices that activate all conditions during training, we enforce the independence of different action formats by enabling only one of them for each training sample. The remaining action conditions will be filled with zeros as unconditional inputs. As demonstrated in Appendix D, this simple constraint prevents the squandering of training cost on action combinations and maximizes the learning efficiency of each individual action mode within the same training steps.
@@ 4.1 Comparisons of Generalization and Fidelity
Automatic Evaluation. Since none of the driving world models are publicly accessible, we compare these methods with their quantitative results on nuScenes. Table 2 reports FID [46] and FVD [113] scores of all methods. Vista exceeds state-of-the-art driving world models with a non-trivial margin.
Human Evaluation. To analyze the generalization of Vista across different datasets, we compare it against three prominent general-purpose video generators trained on web-scale data [5, 131, 142] (see Fig. 5). It is known that automatic metrics like FVD [113] cannot conclusively reveal perceptual quality [3, 6, 31, 128, 134], let alone real-world dynamics. Therefore, we opt for human evaluation for more faithful analysis. Following recent advances [3, 5, 6, 15, 16, 31, 120, 124], we adopt the Two-Alternative Forced Choice protocol. Specifically, participants are presented with a side-by-side video pair and asked to choose the video they deemed better on two orthogonal aspects: visual quality and motion rationality. To avoid potential bias, we crop each video to a fixed aspect ratio, downsample them to the same resolution, and trim the excess frames when Vista generates longer videos than others. We only feed one condition frame to align with other models. To ensure the variety of scenes, we uniformly assemble 60 scenes from four representative datasets: OpenDV-YouTube-val [134], nuScenes [10], Waymo [110], and CODA [77]. These datasets collectively exemplify the intricacy and diversity of real-world driving, e.g., OpenDV-YouTube-val includes geofenced districts, Waymo offers a unique domain compared to our training data, and CODA contains extremely challenging corner cases. We collect a total of 2640 answers from 33 participants. As presented in Fig. 8, Vista outperforms all baselines on both aspects, demonstrating its profound comprehension of the driving dynamics. Further, unlike other models that are only applicable for short-term generation, Vista can accommodate more dynamic priors and produce coherent long-horizon rollouts as shown in Fig. 6.
@@ 4.3 Results of Reward Modeling
4.4 Ablation Study
Dynamic Priors. We visualize the outcomes of applying different orders of dynamic priors in Fig. 11. It shows that dynamic priors play a pivotal role in long-horizon rollouts, where maintaining coherence with respect to historical frames is essential.
Auxiliary Supervisions. To verify the effectiveness of the proposed losses, we devise two additional variants by individually ablating each loss from a variant that incorporates both losses. We compare their effects through Fig. 12, which confirms that the dynamics enhancement loss can promote the learning of real-world dynamics, and the structure preservation loss can reinforce structural details.
@@ Appendix A Discussions
Q8. Broader impact.
@@ B.1 World Models
Intelligent agents should be able to make efficacious decisions even under unseen situations [9, 56, 74, 111, 136, 147]. This requires fundamental knowledge of the world that generalizes to rare cases. As an internal manifestation of such knowledge, a world model predicts plausible futures of the world given potential actions [9, 39, 74, 94, 111, 135]. In principle, it not only predicts how the environment will unfold over time, but also deduces the underlying physical dynamics and agentic behaviors. Such properties can be useful for representation learning [34, 44, 86, 103, 129], model-based reinforcement learning [38, 39, 41, 42, 92, 94, 96], as well as model-predictive control [26, 29, 34, 40, 90, 102, 140]. Recent methods [36, 43, 78, 83, 84] also induce language-based world models from large language models, but are restricted in textual space and struggle with grounding on physics [25, 58].
@@ C.4 Sampling
We generate new videos using the DDIM sampler [107] for 50 steps. The sampling starts with $\sigma_{\text{max}}$ at $700.0$. Unlike SVD that linearly increases the guidance scale, we employ a triangular classifier-free guidance scheme [116] to permit genuine long-horizon rollouts. Specifically, for the $i$-th frame in each $K$ frames to predict, we assign its guidance scale $s(i)$ following:
@@
where $s_{\text{min}}$ and $s_{\text{max}}$ denote the minimum and maximum scales along the frame axis. In our experiments, we define $s_{\text{min}}$ as $1.0$ and $s_{\text{max}}$ as $2.5$. As illustrated in Fig. 14, this technique effectively relieves the saturation drift problem while enhancing details. To improve perceptual continuity, we split the generated latent into clips with an overlap of 3 frames before sending them to the video-aware decoder [5]. The overlapped frames are averaged pixel-wise after decoding.
@@ C.5 Human Evaluation
Recall that we ask the participants to judge side-by-side video pairs from visual quality and motion rationality. To guarantee credible responses, we provide detailed commentary for each aspect of the human evaluation. For visual quality, we let the participants focus on the consistency and harmony of the generated content. For motion rationality, we encourage the participants to pay more attention to the plausibility of the ways that ego-vehicle and other agents move, e.g., whether they are following the traffic rules and exhibiting safe behaviors. For all public models we compared, we use their official configurations for inference and set the prompt as "realistic drive view" if required [131, 142].
@@ C.7 Ablation Studies
For the ablation of loss functions, we train each variant on OpenDV-YouTube [134] for 10K steps at a spatial resolution of 576$\times$1024. All ablations, including the additional ablations in Appendix D, are initialized by loading the pretrained weights of SVD [5] and conducted on 8 A100 GPUs.
@@ Appendix D Additional Experiments
D.1 Additional Results of Reward Modeling
Sensitivity to Hyperparameters. To investigate how the number of denoising steps and the ensemble size influence the performance of the proposed reward function, we repeat the reward estimation procedure in Sec. 4.3 with different hyperparameter settings. We start off by using 5 denoising steps and an ensemble size of 5 for each situation. We then add two variants that double the computational cost by increasing the denoising steps to 10 (our default setting in Appendix C) and increasing the ensemble size to 10 respectively. Following Sec. 4.3, we plot the correlation of the estimated rewards with L2 errors for the three variants in Fig. 13. It shows that increasing the denoising steps can greatly increase the relative contrast of rewards, indicating that denoising step is a more important factor than ensemble size under the same computation budget for reward estimation.
Reward Estimation for Commands. To show that the proposed reward is also practical for other actions, we estimate the rewards of ground truth commands from Waymo and compare them with the rewards of random commands. The results in Table 4 suggest that our reward is also competent for command selection.
D.2 Additional Ablation Studies
Action Independence Constraint. To prove the effectiveness of our learning strategy for action control, we additionally conduct a comparison by removing the action independence constraint proposed in Sec. 3.2. We train two variants on nuScenes [10] at the resolution of 320$\times$576 pixels for 62K steps. The comparison results are shown in Table 4.
Triangular Guidance Scheme. We further compare the introduced classifier-free guidance scheme with the vanilla scheme and the linear scheme [5] to verify its necessity. Fig. 14 shows that our triangular scaling attains the best trade-off between visual quality and saturation preservation.
LoRA Adaptation. To show the necessity of applying LoRA in Sec. 3.2, we train two variants at a resolution of 320$\times$576 pixels for 30K iterations. With the pretrained UNet weights fixed, we let one variant train LoRA and action projection layers in the attention blocks, while the other adjusts new projection layers only. As shown in Fig. 15, adding LoRA is essential for action control learning.
@@ E.1 Generalization Ability
We further demonstrate the strong generalization ability of Vista by deploying it to different scenarios in the wild. The results in Fig. 16 and Fig. 17 show that Vista can make high-fidelity predictions in a very diverse range of scenarios.
@@ E.4 Counterfactual Reasoning Ability
Figure 20: Counterfactual reasoning ability. By imposing actions that violate the traffic rules, we discover that Vista can also predict the consequences of abnormal interventions. In the first example, the ego-vehicle passes over the road boundary and rushes into the bush following our instructions. In the second example, the passing car stops and waits to avoid a collision when we force the ego-vehicle to proceed at the crossroads.

### v5

@@ 2 Preliminary
Despite the high aesthetic quality, SVD lacks several key properties to function as a driving world model. As shown in Sec. 4, the first frame predicted by SVD is not identical to the condition image, making it impractical for autoregressive rollout due to content inconsistency. In addition, SVD struggles with the intricate dynamics of driving scenarios, entailing implausible motions. Moreover, SVD cannot be controlled by any action format. In contrast, we aim to build a generalizable driving world model that predicts high-fidelity futures with realistic dynamics. It ought to be continuously extendable to long horizons and flexibly controllable by multi-modal actions as illustrated in Fig. 2.
@@ 3.2 Phase Two: Learning Versatile Action Controllability
Since the parameters of the camera and vehicle are unavailable for open-world scenarios, it seems impossible to obtain multiple equivalent action conditions simultaneously at inference time. Additionally, it will entail prohibitively expensive training to encompass all possible combinations of action conditions. Hence, unlike common practices that activate all conditions during training, we enforce the independence of different action formats by enabling only one of them for each training sample. The remaining action conditions will be filled with zeros as unconditional inputs. As demonstrated in Appendix D, this simple constraint prevents the squandering of training cost on action combinations and maximizes the learning efficiency of each individual action mode within the same training steps.
@@ 4.1 Comparisons of Generalization and Fidelity
Automatic Evaluation. Since none of the driving world models are publicly accessible, we compare these methods with their quantitative results on nuScenes. We filter 5369 valid samples from the validation set to conduct FID [47] and FVD [115] evaluation. For FID evaluation, we crop and resize the predicted frames to the resolution of 256$\times$448. For FVD evaluation, we use all 25 frames in each video clip and downsample them to 224$\times$224 following LVDM [46]. Table 2 reports the results of all methods. In both metrics, Vista surpasses previous driving world models with a considerable margin.
Human Evaluation. To analyze the generalization of Vista across different datasets, we compare it against three prominent general-purpose video generators trained on web-scale data [5, 133, 144] (see Fig. 5). It is known that automatic metrics like FVD [115] cannot conclusively reveal perceptual quality [3, 6, 32, 130, 136], let alone real-world dynamics. Therefore, we opt for human evaluation for more faithful analysis. Following recent advances [3, 5, 6, 15, 16, 32, 122, 126], we adopt the Two-Alternative Forced Choice protocol. Specifically, participants are presented with a side-by-side video pair and asked to choose the video they deem better on two orthogonal aspects: visual quality and motion rationality. To avoid potential bias, we crop each video to a fixed aspect ratio, downsample them to the same resolution, and trim the excess frames when Vista generates longer videos than others. We only feed one condition frame to align with other models. To ensure the variety of scenes, we uniformly assemble 60 scenes from four representative datasets: OpenDV-YouTube-val [136], nuScenes [10], Waymo [112], and CODA [79]. These datasets collectively exemplify the intricacy and diversity of real-world driving, e.g., OpenDV-YouTube-val includes geofenced districts, Waymo offers a unique domain compared to our training data, and CODA contains extremely challenging corner cases. We collect a total of 2640 answers from 33 participants. As presented in Fig. 8, Vista outperforms all baselines on both aspects, demonstrating its profound comprehension of the driving dynamics. Further, unlike other models that are only applicable for short-term generation, Vista can accommodate more dynamic priors and produce coherent long-horizon rollouts as shown in Fig. 6.
@@ 4.2 Results of Action Controllability
We also introduce a new metric named Trajectory Difference to assess control consistency. Following GenAD [136], we train an inverse dynamics model (IDM) that estimates the corresponding trajectory from a video clip. An illustration of IDM is shown in Fig. 13. We then send Vista’s prediction to the IDM and calculate the L2 difference between the ground truth trajectory and the estimated trajectory. The differences are measured over 2 seconds. The lower the trajectory difference, the stronger the control consistency Vista exhibits. We conduct the experiments on nuScenes and Waymo. For each dataset, we collect a subset that contains 537 samples. As reported in Table 3, Vista can be effectively controlled by different modalities of actions, resulting in more consistent motions to the ground truth.
@@ 4.3 Results of Reward Modeling
4.4 Ablation Study
Dynamic Priors. Fig. 11 visualizes the outcomes of using different orders of dynamic priors. The order of priors corresponds to the number of condition frames. It shows that dynamic priors play a pivotal role in long-horizon rollouts, where the coherence with respect to historical frames is essential.
Figure 13: An illustration of the IDM experiments in Table 3.
To further demonstrate the efficacy of dynamic priors, we conduct a quantitative evaluation in Table 3. Specifically, we use the IDM in Sec. 4.2 to infer the trajectories of the predicted videos with different orders of dynamic priors. The diminishing differences in trajectory suggest that introducing more priors can effectively improve the consistency between prediction and ground truth.
Auxiliary Supervisions. To verify the effectiveness of the two losses proposed in Sec. 3.1, we devise two additional variants by individually ablating each loss from a variant that incorporates both losses. We qualitatively compare their effects through Fig. 12, which confirms that the dynamics enhancement loss can promote the learning of real-world dynamics, and the structure preservation loss can reinforce the prediction of structural details.
Table 3: Impacts of different action conditions and dynamic priors. By applying action conditions and dynamic priors, Vista can predict motion that is more consistent compared to the ground truth.
Dataset Condition Average Trajectory Difference $\downarrow$
with 1 prior with 2 priors with 3 priors
nuScenes GT video 0.379 0.379 0.379
action-free 3.785 2.597 1.820
+ goal point 2.869 2.192 1.585
+ command 3.129 2.403 1.593
+ angle & speed 1.562 1.123 0.832
+ trajectory 1.559 1.148 0.835
Waymo GT video 0.893 0.893 0.893
action-free 3.646 2.901 2.052
+ command 3.160 2.561 1.902
+ trajectory 1.187 1.147 1.140
@@ Appendix A Discussions
Q8. Why not expand the Vista framework to surround-view generation?
It is true that supporting surround-view generation would further help driving. Existing works [127, 145] have explored the surround-view settings on nuScenes [10]. However, in this paper, we focus on the front-view setting for three main reasons: (1) The front view setting allows leveraging diverse data sources [54, 136]. Conversely, the distinctions in multi-view videos from various datasets, such as different numbers of cameras, hinder unified modeling and data scaling. (2) Models that focus on the front view can be seamlessly applied to different datasets without adaptation [107], broadening their applicability across datasets. (3) Though incomplete, the front view often contains most of the information vital for driving. As demonstrated in NAVSIM [23], using the front view alone results in only a $1.1\%$ performance downtick in collision rate compared to using five surround-view cameras.
Q9. Broader impact.
@@ B.1 World Models
Intelligent agents should be able to make efficacious decisions even under unseen situations [9, 57, 76, 113, 138, 149]. This requires fundamental knowledge of the world that generalizes to rare cases. As an internal manifestation of such knowledge, a world model predicts plausible futures of the world given potential actions [9, 40, 69, 76, 96, 113, 137]. In principle, it not only predicts how the environment will unfold over time, but also deduces the underlying physical dynamics and agentic behaviors. Such properties can be useful for representation learning [35, 45, 88, 105, 131], model-based reinforcement learning [39, 40, 42, 43, 94, 96, 98], as well as model-predictive control [27, 30, 35, 41, 92, 104, 142]. Recent methods [37, 44, 80, 85, 86] also induce language-based world models from large language models, but are restricted in textual space and struggle with grounding on physics [26, 59].
@@ C.4 Sampling
We generate future videos using the DDIM sampler [109] for 50 steps. The sampling starts with $\sigma_{\text{max}}$ at $700.0$. Since our model predicts long-term futures in an autoregressive manner, the issue of over-saturation caused by the standard classifier-free guidance will accumulate rapidly. Therefore, unlike SVD that linearly increases the guidance scale, we employ a triangular classifier-free guidance scheme [118] to permit genuine long-horizon rollouts. Specifically, for the $i$-th frame in each $K$ frames to predict, we assign its guidance scale $s(i)$ following:
@@
where $s_{\text{min}}$ and $s_{\text{max}}$ indicate the minimum and maximum guidance scales along the temporal axis. In our experiments, we define $s_{\text{min}}$ as $1.0$ and $s_{\text{max}}$ as $2.5$. This triangle scheme assigns moderate guidance scales to the frames that will be used as conditions in the next prediction round. Due to sufficient temporal interaction, the quality of intermediate frames can also propagate to the frames that have lower guidance scales. As illustrated in Fig. 15, this technique adeptly mitigates the saturation drift problem while enhancing details. To improve perceptual continuity, we split the generated latent into clips with an overlap of 3 frames before sending them to the video-aware decoder [5]. After decoding, the overlapped frames are averaged pixel-wise.
@@ C.5 Human Evaluation
Recall that we ask the participants to judge side-by-side video pairs from visual quality and motion rationality. To guarantee credible responses, we provide detailed commentary for each aspect of the human evaluation. For visual quality, we let the participants focus on the consistency and harmony of the generated content. For motion rationality, we encourage the participants to pay more attention to the plausibility of the ways that ego-vehicle and other agents move, e.g., whether they are following the traffic rules and exhibiting safe behaviors. For all public models we compared, we use the official checkpoints and configurations for inference without finetuning. For the models that require textual inputs [133, 144], we set the prompt as "realistic drive view".
@@ C.7 Ablation Studies
For the ablation of loss functions, we train each variant on OpenDV-YouTube [136] for 10K steps at a spatial resolution of 576$\times$1024. All ablations, including the additional ablations in Appendix D, are initialized by loading the pretrained checkpoint of SVD [5] and conducted with 8 A100 GPUs.
@@ Appendix D Additional Experiments
D.1 Parameter Sensitivity of Reward Estimation
To investigate how the number of denoising steps and the ensemble size influence the performance of the proposed reward function, we repeat the reward estimation procedure in Sec. 4.3 with different hyperparameter settings. We start off by using 5 denoising steps and an ensemble size of 5 for each sample. We then test two variants that double the computational cost by increasing the denoising steps to 10 (our default setup in Appendix C) and increasing the ensemble size to 10 respectively. Following Sec. 4.3, we plot the correlation of the estimated rewards with L2 errors for the three variants in Fig. 14. The results show that increasing the number of denoising steps can greatly enlarge the relative contrast of rewards, suggesting that denoising step is a more important factor than ensemble size under the same computation budget for reward estimation.
D.2 Reward Estimation for Commands
To demonstrate that the proposed reward function is also applicable for other actions, we estimate the rewards of ground truth commands from Waymo and compare them with the rewards of random commands. The results in Table 5 suggest that our reward is also competent for command selection.
D.3 Action Independence Constraint
To prove the efficacy of our learning strategy for action control, we conduct a comparison by removing the action independence constraint proposed in Sec. 3.2. We train two variants on nuScenes [10] at the resolution of 320$\times$576 pixels for 62K steps. The comparison results are presented in Table 5.
@@ D.3 Action Independence Constraint
D.4 Triangular Guidance Scheme
We further compare the introduced classifier-free guidance scheme with the vanilla scheme and the linear scheme [5] to verify its necessity. Fig. 15 shows that our triangular scaling attains the best trade-off between visual quality and saturation preservation.
D.5 LoRA Adaptation
To show the necessity of applying LoRA in Sec. 3.2, we train two variants at a resolution of 320$\times$576 pixels for 30K iterations. With the pretrained UNet weights fixed, we let one variant train LoRA and action projection layers in the attention blocks, while the other adjusts new projection layers only. As shown in Fig. 16, adding LoRA is essential for action control learning.
D.6 Action Control Consistency
In Table 6, we report the complete FVD scores of Fig. 8, which further validates the effectiveness of all kinds of action controls. Note that since our "stop" subset consists of samples where the final displacements are within 2 meters, the goal points typically do not appear in these videos. Hence, for the experiment that uses goal point as action condition on the "stop" subset, most samples are generated in the same way as the action-free mode.
D.7 Human Evaluation with GenAD
To our best knowledge, there is not a driving-specific world model publicly available so far, making it hard to conduct qualitative human evaluation. Hence, we mainly compare Vista against the existing methods with the officially reported FID and FVD scores in Table 2.
Figure 17: Perceptual disparity between GenAD and Vista.
To demonstrate the considerable improvements in visual quality and motion rationality, we conduct an extra human evaluation with the state-of-the-art GenAD model [136]. Since GenAD processes a 4-second video each time, we perform autoregressive prediction to extend Vista ’s output to 5 seconds and trim the last second to align with GenAD’s duration. To avoid any bias caused by resolution and frequency, we downsample the outputs of Vista (576$\times$1024 resolution at 10 Hz) to 256$\times$448 resolution at 2 Hz. The evaluation follows the same procedure specified in Sec. 4.1.
We collect 25 diverse samples from the unseen OpenDV-YouTube-val set and invite 20 volunteers for evaluation. We ask the volunteers to choose the video they deem better. As a result, Vista is preferred in $94.4\%$ and $94.8\%$ of the time on visual quality and motion rationality respectively. This attests that Vista, in spite of experiencing a large perceptual reduction due to downsampling, still exhibits a significant advantage over GenAD in generation quality. We also compare the predictions of GenAD and Vista in Fig. 17, showing the superiority of Vista in resolution and fidelity.
Table 6: Complete FVD scores of different action categories. We obtain the FVD scores on four subsets divided by command categories. All types of action controls are effective across all categories.
Dataset Condition Subset FVD $\downarrow$
forth right left stop average
nuScenes action-free 135.6 405.6 513.8 414.1 367.2
+ goal point 122.4 315.6 439.6 413.5 322.7
+ command 122.2 299.7 485.6 261.6 292.2
+ angle & speed 122.8 285.6 397.8 114.1 230.0
+ trajectory 125.2 229.2 357.7 118.5 207.6
Waymo action-free 145.9 407.6 529.9 164.1 311.8
+ command 122.5 331.5 496.9 143.9 273.7
+ trajectory 126.3 285.5 527.6 136.5 268.9
@@ E.1 Generalization Ability
We further demonstrate the strong generalization ability of Vista by deploying it to different scenarios in the wild. The results in Fig. 18 and Fig. 19 illustrate that Vista can make high-fidelity predictions in a very diverse range of scenarios.
@@ E.4 Counterfactual Reasoning Ability
Figure 22: Counterfactual reasoning ability. By imposing actions that violate the traffic rules, we discover that Vista can also predict the consequences of abnormal interventions. In the first example, the ego-vehicle passes over the road boundary and rushes into the bush following our instructions. In the second example, the passing car stops and waits to avoid a collision when we force the ego-vehicle to proceed at the crossroads. This showcases Vista’s potential for facilitating closed-loop simulation.